OpenStudy (anonymous):

Janz is putting turf on a triangular portion of a garden around a circular fountain. If the turf costs $10 a square meter, how much will it cost Janz? Round to the nearest cent.

OpenStudy (anonymous):



OpenStudy (dumbsearch2):

work out area of triangle, formula is half of the base length times the height of the triandle. then work out the area of the circle, formula is pi times the radius of the circle squared. then subtract the circle are from the triangle area and you will have the area you want. multiply final by $10 to get the total cost. make sure you change all your lengths to metres before you work out the areas.

OpenStudy (dumbsearch2):

Does that help? :)

OpenStudy (anonymous):

Perfect!

OpenStudy (dumbsearch2):

:))

OpenStudy (anonymous):

Okay not perfect I did something wrong, my options are $ 5611.84 $ 4480.87 $ 5432.84 $ 6480.00 but I got 534.96 .-.

OpenStudy (dumbsearch2):

alright... give me a minute or many minutes please

OpenStudy (dumbsearch2):

to figure this one out...

OpenStudy (anonymous):

Okay, I appreciate it.

OpenStudy (dumbsearch2):

give me a little while, almost done. sry for the wait.

OpenStudy (anonymous):

No worries

OpenStudy (dumbsearch2):

Thank you. Your quite nice.

OpenStudy (dumbsearch2):

Try bumping your question occasionally to give it the maximum exposure, so everyone sees it, so everyone can try to help. See?

OpenStudy (dumbsearch2):

click the "bump" button near your "open" question. see? :D

OpenStudy (anonymous):

I can't bump it yet -______________-

OpenStudy (dumbsearch2):

bump this question now, Im still confused, trying to work this out...

OpenStudy (anonymous):

it wont let my bump it says bump in 1400:23 D:

OpenStudy (dumbsearch2):

wa???

OpenStudy (anonymous):

1391:00 now.......

OpenStudy (dumbsearch2):

Tell me if this helps, but I found: Area of equilateral triangle = 1/2 * s² * √3/2 = s²√3/4 Area of triangle = 36²√3/4 = 324√3 Area of circular fountain = π * 6² = 36π Area of turf = (324√3 − 36π) sq m Cost = (324√3 − 36π) sq m * $10/sq m = $4480.87

OpenStudy (dumbsearch2):

This works?